# Altermagnetic superconducting diode effect

### v1

@@ III.1 General formalism
$\mathcal{H}_{I}=-g\sum_{\boldsymbol{q}}C^{\dagger}_{\boldsymbol{q}}C_{\boldsymbol{q}},\quad C_{\boldsymbol{q}}=\sum_{\boldsymbol{k}}c_{\boldsymbol{k}+\boldsymbol{q}/2}is_{y}c_{-\boldsymbol{k}+\boldsymbol{q}/2},$ (8)
@@
$\begin{split}\mathcal{H}_{\text{MF}}=&\sum_{\boldsymbol{k}}c_{\boldsymbol{k}}^{\dagger}h_{\boldsymbol{k}}c_{\boldsymbol{k}}+\sum_{\boldsymbol{k},\boldsymbol{q}}\left[\Delta_{\boldsymbol{q}}c^{\dagger}_{\boldsymbol{k}+\boldsymbol{q}/2}is_{y}c^{\dagger}_{-\boldsymbol{k}+\boldsymbol{q}/2}+\text{H.c.}\right]\\ &+\sum_{\boldsymbol{q}}\frac{|\Delta_{\boldsymbol{q}}|^{2}}{g},\quad h_{\boldsymbol{k}}=s_{0}\epsilon_{\boldsymbol{k}}+\boldsymbol{g}_{\boldsymbol{k}}\cdot\boldsymbol{s}+\boldsymbol{N}_{\boldsymbol{k}}\cdot\boldsymbol{s}.\end{split}$ (9)
@@
$\mathcal{H}_{\text{MF}}(\boldsymbol{q})=\sum_{\boldsymbol{k}}\Psi_{\boldsymbol{k}}^{\dagger}H_{\text{BdG}}\Psi_{\boldsymbol{k}}+\frac{|\Delta_{\boldsymbol{q}}|^{2}}{g}.$ (11)
@@ III.3 Generic zeros in $C_{6h}$
Figure 3: SDE in the point group $C_{6h}$ at $E_{z}\neq 0$. (a) Particle-particle bubble $\Gamma(\boldsymbol{q})$ in the Brillouin zone; (b) shows the critical current $J_{c}(\hat{\boldsymbol{n}})$ as a function of direction $\hat{\boldsymbol{n}}$; (c) displays the angle-resolved SDE efficiency $\eta(\hat{\boldsymbol{n}})$. Due to the reduced symmetry compared to $D_{6h}$ [cf. Fig. 2b(iv)], the zeros (and also maxima) of $\eta$ are rotated away from high-symmetry directions. We choose $N_{0}=-0.21,\alpha=0.4,t^{\prime}=0.05t$, $g^{-1}=0.5\Gamma(\boldsymbol{q}=0)$, $\varphi=-0.25\pi$, $\varphi^{\prime}=0$.
@@ IV Conclusion
In this work, we have studied under which conditions a conventional two-dimensional superconducting phase can exhibit a SDE, $\eta_{s}\neq 0$ in Eq. (2), if the only source of broken time-reversal symmetry is an AM order parameter. To approach this question systematically, we started by listing all AM order parameters $\boldsymbol{N}_{\boldsymbol{k}}$, coupling to the electrons as given by Eq. (5), for all possible 2D crystalline point groups, see Table 1. For each of them, we checked whether there are still residual symmetries, such as inversion symmetry or two-fold rotation along a direction perpendicular to the plane of the system, that would enforce the critical currents along opposite directions to be the same, $J_{c}(\hat{\boldsymbol{n}})=J_{c}(-\hat{\boldsymbol{n}})$. We found that there are indeed several point groups and associated AM order parameters ($C_{6h}$ with AM in $B_{g}$; $D_{6h}$ with AM in $B_{1g}$ and $B_{2g}$; $D_{3d}$ with $A_{1g}$; $D_{3}$ with AM in $A_{1}$; $D_{6}$ with $B_{1}$ and $B_{2}$) that give rise to an SDE, as we have also checked by explicit model calculations. Since, by design, the total magnetic moment vanishes when an AM is the only source of broken time-reversal symmetry, this demonstrates that the SDE does not require finite net magnetization.
@@ Appendix B Details of the models
Figure 5: Multi-component momentum dependence of (a) the altermagnet term $\boldsymbol{N}_{\boldsymbol{k}}$; (b) inversion-symmetry breaking term $\boldsymbol{g}_{\boldsymbol{k}}^{I}$ and (c) shows the spin-orbit coupling term $\boldsymbol{g}_{\boldsymbol{k}}^{II}$ induced by $E_{z}$. For point groups $D_{6h}$ and $C_{6h}$, $\nu=0$ and $\beta=0$, while for $D_{3}$ $\nu=1$ and $\beta\neq 0$.
$\boldsymbol{N}_{\boldsymbol{k}}=N_{0}\left(\frac{-8}{3}(\cos{k_{x}}-\cos{\frac{k_{x}}{2}}\cos{\frac{\sqrt{3}k_{y}}{2}}),\frac{-8\sin{\frac{k_{x}}{2}}\sin{\frac{\sqrt{3}k_{y}}{2}}}{\sqrt{3}},64\nu\frac{(\sin{\frac{5k_{x}}{2}}\sin{\frac{3\sqrt{3}k_{y}}{2}}+\sin{\frac{k_{x}}{2}}\sin{\frac{\sqrt{3}k_{y}}{2}}-\sin{2k_{x}}\sin{\sqrt{3}k_{y}})}{3\sqrt{3}}\right)$ (22)
@@
$\boldsymbol{g}^{\text{II}}_{\boldsymbol{k}}=\alpha\left(\frac{2}{\sqrt{3}}\cos{\frac{k_{x}}{2}}\sin{\frac{\sqrt{3}k_{y}}{2}}\right.,\left.{}\frac{-2}{3}(2\cos{\frac{k_{x}}{2}}+\cos{\frac{\sqrt{3}k_{y}}{2}})\sin{\frac{k_{x}}{2}},16\nu\left(-\cos{\frac{k_{x}}{2}}+\cos{\frac{\sqrt{3}k_{y}}{2}}\right)\sin{\frac{k_{x}}{2}}\right).$ (24)
These are the functions with the least number of nodes/nodal lines that satisfy the required symmetry constraints while being periodic on the Brillouin zone. As explained in the main text, $\nu=0$ for point groups $D_{6h}$ and $C_{6h}$ and therefore Eq. (22) reduces to Eq. (13) and Eq. (15) respectively when expanded for small $\boldsymbol{k}$. As the point group $D_{3}$ permits the AM order parameter to also have a third component, it is given fully by Eq. (22) with $\nu=1$ resulting in Eq. (16) for $\boldsymbol{k}\rightarrow 0$. On the same token, the leading basis functions for the electric-field induced spin-orbit coupling $\boldsymbol{g}^{\text{II}}_{\boldsymbol{k}}$ are given by Eq. (24), with $\nu=0$ for $\mathcal{G}=D_{6h}$ and $C_{6h}$ and $\nu=1$ for $\mathcal{G}=D_{3}$. Finally, the leading basis function for $\boldsymbol{g}^{\text{I}}_{\boldsymbol{k}}\sim\beta(k_{x},k_{y},k_{y}(3k_{x}^{2}-k_{y}^{2}))$ is given by Eq. (23).

### v2

@@ III.1 General formalism
$\mathcal{H}_{I}=-g\sum_{\boldsymbol{q}}C^{\dagger}_{\boldsymbol{q}}C^{\phantom{\dagger}}_{\boldsymbol{q}},\quad C^{\phantom{\dagger}}_{\boldsymbol{q}}=\sum_{\boldsymbol{k}}c_{\boldsymbol{k}+\boldsymbol{q}/2}is_{y}c_{-\boldsymbol{k}+\boldsymbol{q}/2},$ (8)
@@
$\begin{split}\mathcal{H}_{\text{MF}}=&\sum_{\boldsymbol{k}}c_{\boldsymbol{k}}^{\dagger}h_{\boldsymbol{k}}c_{\boldsymbol{k}}^{\phantom{\dagger}}+\sum_{\boldsymbol{k},\boldsymbol{q}}\left[\Delta_{\boldsymbol{q}}c^{\dagger}_{\boldsymbol{k}+\boldsymbol{q}/2}is_{y}c^{\dagger}_{-\boldsymbol{k}+\boldsymbol{q}/2}+\text{H.c.}\right]\\ &+\sum_{\boldsymbol{q}}\frac{|\Delta_{\boldsymbol{q}}|^{2}}{g},\quad h_{\boldsymbol{k}}=s_{0}\epsilon_{\boldsymbol{k}}+\boldsymbol{g}_{\boldsymbol{k}}\cdot\boldsymbol{s}+\boldsymbol{N}_{\boldsymbol{k}}\cdot\boldsymbol{s}.\end{split}$ (9)
@@
$\mathcal{H}_{\text{MF}}(\boldsymbol{q})=\sum_{\boldsymbol{k}}\Psi_{\boldsymbol{k}}^{\dagger}H_{\text{BdG}}\Psi_{\boldsymbol{k}}^{\phantom{\dagger}}+\frac{|\Delta_{\boldsymbol{q}}|^{2}}{g}.$ (11)
@@ III.3 Generic zeros in $C_{6h}$
We further present in Fig. 3(d) the diode efficiency $\eta$ for both the magnetic field-induced SDE and the altermagnetic SDE. To treat them on equal footing, we use the same model and set $\boldsymbol{B}^{2}=N_{0}^{2}=\sum_{\boldsymbol{k}\in\text{BZ}}\boldsymbol{N}_{\boldsymbol{k}}^{2}$. Employing the same parameters as in the other parts of Fig. 3, we find that the efficiency is significantly larger in the AM case; this demonstrates that the AM SDE is not a subleading and, thus generally, weaker effect compared to the magnetic-field-induced one, contrary to what one might have expected since the average magnetization is required to vanish for an AM. However, we caution that both magnetic fields and AMs can yield sizeable SDE with relative magnitudes depending largely on the parameters chosen.
Figure 3: SDE in the point group $C_{6h}$ at $E_{z}\neq 0$. (a) Particle-particle bubble $\Gamma(\boldsymbol{q})$ in the Brillouin zone; (b) shows the critical current $J_{c}(\hat{\boldsymbol{n}})$ as a function of direction $\hat{\boldsymbol{n}}$ for $N_{0}=-0.21$; (c) displays the angle-resolved SDE efficiency $\eta(\hat{\boldsymbol{n}})$. Due to the reduced symmetry compared to $D_{6h}$ [cf. Fig. 2b(iv)], the zeros (and also maxima) of $\eta$ are rotated away from high-symmetry directions; (d) SDE efficiency $\eta$ as a function of magnetic field and altermagnetic order parameter $N_{0}$ shown in blue and red respectively. The magnetic field is chosen along the $x$ direction. In this figure, we used $\alpha=0.4,t^{\prime}=0.05t$, $g^{-1}=0.5\Gamma(\boldsymbol{q}=0)$, $\varphi=-0.25\pi$, $\varphi^{\prime}=0$.
@@ IV Conclusion
In this work, we have studied under which conditions a conventional two-dimensional superconducting phase can exhibit a SDE, $\eta_{s}\neq 0$ in Eq. (2), if the only source of broken time-reversal symmetry is an AM order parameter. To approach this question systematically, we started by listing all AM order parameters $\boldsymbol{N}_{\boldsymbol{k}}$, coupling to the electrons as given by Eq. (5), for all possible 2D crystalline point groups, see Table 1. For each of them, we checked whether there are still residual symmetries, such as inversion symmetry or two-fold rotation along a direction perpendicular to the plane of the system, that would enforce the critical currents along opposite directions to be the same, $J_{c}(\hat{\boldsymbol{n}})=J_{c}(-\hat{\boldsymbol{n}})$. We found that there are indeed several point groups and associated AM order parameters ($C_{6h}$ with AM in $B_{g}$; $D_{6h}$ with AM in $B_{1g}$ and $B_{2g}$; $D_{3d}$ with $A_{1g}$; $D_{3}$ with AM in $A_{1}$; $D_{6}$ with $B_{1}$ and $B_{2}$) that give rise to an SDE, as we have also checked by explicit model calculations. Since, by design, the total magnetic moment vanishes when an AM is the only source of broken time-reversal symmetry, this demonstrates that the SDE does not require finite net magnetization. A direct comparison with the magnetic-field-induced SDE reveals that the altermagntic SDE is not a subleading effect; in fact, depending on parameters, the altermagnetic efficiency can even be substantially larger when compared in the same model with the magnetic-field scenario.
@@ Appendix B Details of the models
Figure 5: Multi-component momentum dependence of (a) the altermagnet term $\boldsymbol{N}_{\boldsymbol{k}}$; (b) inversion-symmetry breaking term $\boldsymbol{g}_{\boldsymbol{k}}^{I}$ and (c) shows the spin-orbit coupling term $\boldsymbol{g}_{\boldsymbol{k}}^{II}$ induced by $E_{z}$. For point groups $D_{6h}$ and $C_{6h}$, $\tau=0$, $\tilde{\tau}=0$ and $\beta=0$, while for $D_{3}$, $\tau\neq 0$, $\tilde{\tau}\neq 0$ and $\beta\neq 0$.
$\boldsymbol{N}_{\boldsymbol{k}}=N_{0}\left(\frac{-8}{3}(\cos{k_{x}}-\cos{\frac{k_{x}}{2}}\cos{\frac{\sqrt{3}k_{y}}{2}}),\frac{-8\sin{\frac{k_{x}}{2}}\sin{\frac{\sqrt{3}k_{y}}{2}}}{\sqrt{3}},64\tau\frac{(\sin{\frac{5k_{x}}{2}}\sin{\frac{3\sqrt{3}k_{y}}{2}}+\sin{\frac{k_{x}}{2}}\sin{\frac{\sqrt{3}k_{y}}{2}}-\sin{2k_{x}}\sin{\sqrt{3}k_{y}})}{3\sqrt{3}}\right)$ (22)
@@
$\boldsymbol{g}^{\text{II}}_{\boldsymbol{k}}=\alpha\left(\frac{2}{\sqrt{3}}\cos{\frac{k_{x}}{2}}\sin{\frac{\sqrt{3}k_{y}}{2}}\right.,\left.{}\frac{-2}{3}(2\cos{\frac{k_{x}}{2}}+\cos{\frac{\sqrt{3}k_{y}}{2}})\sin{\frac{k_{x}}{2}},16\tilde{\tau}\left(-\cos{\frac{k_{x}}{2}}+\cos{\frac{\sqrt{3}k_{y}}{2}}\right)\sin{\frac{k_{x}}{2}}\right).$ (24)
These are the functions with the least number of nodes/nodal lines that satisfy the required symmetry constraints while being periodic on the Brillouin zone. As explained in the main text, $\tau=0$ for point groups $D_{6h}$ and $C_{6h}$ and therefore Eq. (22) reduces to Eq. (13) and Eq. (15), respectively, when expanded for small $\boldsymbol{k}$. As the point group $D_{3}$ permits the AM order parameter to also have a third component, it is given by Eq. (22) with $\tau\neq 0$ resulting in Eq. (16) for $\boldsymbol{k}\rightarrow 0$. Similarly, the leading basis functions for the electric-field induced spin-orbit coupling $\boldsymbol{g}^{\text{II}}_{\boldsymbol{k}}$ are given by Eq. (24), with $\tilde{\tau}=0$ for $\mathcal{G}=D_{6h}$ and $C_{6h}$; meanwhile $\tilde{\tau}\neq 0$ for $\mathcal{G}=D_{3}$. Finally, the leading spin-orbit vector at zero electric field for the point group $D_{3}$ is given by $\boldsymbol{g}^{\text{I}}_{\boldsymbol{k}}$ in Eq. (23).